# Learning Non-Ideal Single Vortex Flows Using the Differentiable Vortex Particle Method

### v2

@@ 6 Conclusion
In this study, we extend the capabilities of the DVPM to non-ideal conditions, encompassing viscous flows and non-conservative body forces. Furthermore, we employ the Lamb-Oseen vortex as a more rigorous benchmark for model evaluation, as it provides an exact solution to the NS equations. However, our current framework exhibits limitations in accurately predicting multiple-vortex systems, primarily due to insufficient incorporation of the physical laws governing vortex interactions. Beyond this limitation, future work will address additional challenges, including flow transition, turbulent, and three-dimensional flow configurations.
@@ Appendix A Implementation detail
In this section, we describe the implementation details of our proposed method.
The networks $N_{\boldsymbol{x}}$, $N_{\omega}$, and $N_{\delta^{2}}$ are implemented through a sequence of 3 residual blocks with progressively increasing widths of $[64,128,256]$. The network $N_{K}$ comprises four fully connected layers, each containing 40 neurons. Additional hyperparameters align with those established in previous research [2]. Prior to the primary training phase, $N_{\boldsymbol{x}}$ undergoes a pretraining process for 10,000 iterations with two distinct objectives: (1) ensuring that for all $t\in\left[0,t_{E}\right]$, $N_{\boldsymbol{x}}(t)=\left[\left(\boldsymbol{x}_{1}\right)_{t},\ldots,\left(\boldsymbol{x}_{n}\right)_{t}\right]$ precisely corresponds to the centers of an $n\times n$ grid, and (2) establishing that for all $t\in\left[0,t_{E}\right]$, $\frac{dN_{\boldsymbol{x}}}{dt}=0$, thereby initializing the particles in a stationary state.
@@
Appendix B Single Lamb-Oseen vortex with Coriolis force
[…]
Given that a Lamb-Oseen vortex under Coriolis force lacks an exact analytical solution, we employ the CFD software OpenFOAM for numerical simulation. We utilize a circular computational domain for this investigation to mitigate boundary condition effects on both internal and external flows. Figure 9 illustrates the computational domain configuration and spatial discretization methodology implemented in our CFD simulation. We partition the computational domain into five distinct blocks to enhance mesh quality. The grid point distribution for the upper-left block is explicitly indicated in the figure, while the remaining blocks maintain symmetrical grid point distributions concerning the coordinate origin. The sampling interval and boundary conditions employed in this analysis are identical to those described in Section 5.2.
Figure LABEL:lamb_oseen_vortex_Re_1000_comparison illustrates the comparative evolution of the Lamb-Oseen vortex at $t=1.2\text{s}$ ($Re=1000$) with and without the influence of the Coriolis force. The results demonstrate that when the rotational direction of the computational domain aligns with that of the Lamb-Oseen vortex (counter-clockwise rotation), the Coriolis force enhances the vortex structure. Conversely, the vortex is weakened when the rotational directions oppose each other. This observed phenomenon is consistent with the theoretical predictions derived from Equation 1.
Figure 11 presents a comparative analysis of the prediction results obtained from our proposed model and several baseline models. The results demonstrate that our model accurately predicts the single vortex flow dynamics under viscous effects and Coriolis forces. While the CNN-based model successfully captures most of the vortex flow characteristics, it exhibits notable limitations in accurately resolving the core region. In contrast, the PINN-based model yields significantly inferior results compared to the other two approaches, with substantial deviations from the expected flow patterns.
Appendix C Training loss curve of PINN for modeling multiple Lamb–Oseen vortices
Figure 12 illustrates the training loss curve of the PINN for modeling multiple Lamb–Oseen vortices, as described in Section 5.2. The results indicate that the loss function has reached a plateau, and further reduction appears unattainable. Despite this convergence in the training process, the results presented in Figure 7 remain unsatisfactory.
Figure 12: Training loss curve of PINN for modeling multiple Lamb–Oseen vortices.

### v3

@@ 4 Conclusion
In this study, we extend the capabilities of the DVPM to non-ideal conditions, specifically encompassing viscous flows and non-conservative body forces. Furthermore, we utilize the Lamb-Oseen vortex as a more rigorous benchmark for model evaluation, given that it provides an exact solution to the NS equations.
In detail, we employed the Lamb-Oseen vortex at Reynolds numbers of 10, 100, and 1000 as training and test datasets to validate our proposed method. We conducted comprehensive comparisons with PINN-based and CNN-based approaches, as well as the original DVPM. The results demonstrate that our method outperforms these baseline methods across nearly all evaluation metrics. Furthermore, we tested our approach using the Lamb-Oseen vortex with Coriolis force, which also yielded satisfactory results.
Future work will focus on integrating additional physical functionalities derived from traditional VPM research to simulate more complex real engineering flow problems, thereby combining the advantages of both the traditional VPM and deep learning.
Acknowledgments and Disclosure of Funding
This research was supported by National Natural Science Foundation of China [No. 52406057], Research Grants Council of Hong Kong [No. CityU 21206123] and National Natural Science Foundation of Guangdong Province [Grant No. 2514050003981].
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used GPT-4 in order to improve readability and language. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Appendix A Implementation detail
In this section, we describe the implementation details of our proposed method. The networks $N_{\boldsymbol{x}}$, $N_{\omega}$, and $N_{\delta^{2}}$ are implemented through a sequence of 3 residual blocks with progressively increasing widths of $[64,128,256]$. The network $N_{K}$ comprises four fully connected layers, each containing 40 neurons. Additional hyperparameters align with those established in previous research [15]. Prior to the primary training phase, $N_{\boldsymbol{x}}$ undergoes a pretraining process for 10,000 iterations with two distinct objectives: (1) ensuring that for all $t\in\left[0,t_{E}\right]$, $N_{\boldsymbol{x}}(t)=\left[\left(\boldsymbol{x}_{1}\right)_{t},\ldots,\left(\boldsymbol{x}_{n}\right)_{t}\right]$ precisely corresponds to the centers of an $n\times n$ grid, and (2) establishing that for all $t\in\left[0,t_{E}\right]$, $\frac{dN_{\boldsymbol{x}}}{dt}=0$, thereby initializing the particles in a stationary state.